OpenStudy (anonymous):

Multiply. 1.) 7i (3-7i) 2.) (4+3i) (3-3i)

OpenStudy (anonymous):

1.) 7i (3-7i)=7i *3 - 7i * 7i = 21i - 49*i^2=49+21i ; 2.) (4+3i) (3-3i) = 4*3+4*(-3i)+3i*3+3i*(-3i)=12-12i+9i-9*i^2=12-3i+9=21-3i

OpenStudy (anonymous):

is that by rationalizing ?

OpenStudy (anonymous):

ah, rationalizing is a process to turn the denominator which is probably not a rational to a national.

OpenStudy (anonymous):

rational,...

OpenStudy (anonymous):

rational,...

OpenStudy (anonymous):

so if i wanted to rationalize 10/1+3i ?

OpenStudy (anonymous):

you mean 10/(1+3i)?? you'd better use the brackets to avoid misunderstanding

OpenStudy (anonymous):

and same process if i want to add or subtract ?

OpenStudy (anonymous):

yes 10/(1+3i), sorry i've never done math online

OpenStudy (anonymous):

you understand the two methods: traditional one, and rationalization??

OpenStudy (anonymous):

rationalization is longer ?